A copper wire has a resistance of 85Ω at 150C. Calculate its resistance at 900C. Assuming the temperature coefficient resistance of copper to be 0.0043/0C,


1
Expert's answer
2021-09-02T10:42:27-0400

At 15°C:


"R_1=R_0(1+\alpha T_1),\\\space\\
R_0=\frac{R_1}{1+\alpha T_1}."

Find the resistance at 90°C:

"R_2=R_0(1+\alpha T_2),\\\space\\
R_2=\frac{R_1}{1+\alpha T_1}(1+\alpha T_2)=110\Omega."